OpenStudy (anonymous):

At one factory, the ratio of defective light bulbs produced to total light bulbs produced is about 3:500. How many light bulbs are expected to be defective when 12,000 are produced?

OpenStudy (anonymous):

\[\frac{3}{500}=\frac{x}{12,000}\iff x=\frac{3\times 12,000}{500}\]